if a certain amounts to rs.44,000 at 5% p.a over 2 years, how much will it amount to be ub 3 years at the same rate? Answer by mananth(16949) 
 (44000-p)=pnr/100
100(44000-p)=pnr
r=5
n=2
4400000-100p=p*5*2
4400000-100p=10p
4400000=100p+10p
4400000=110p
4400000/110=p
=40,000
P=Rs.40,000
Interest for 2 years = Rs.4000
For 1 year it will be Rs.2000
So for 3 years it will be Rs. 6000
m.ananth@hotmail.ca